# Bundle Neural Networks for message diffusion on graphs

### v1

@@ Appendix F Experiment details
For peptides-func and peptides-struct we use a fixed parameter budget of $\sim 500k$ as in Dwivedi et al. [32]. We fix hyper-parameters to be the best GCN hyper-parameters from Tönshoff et al. [36], and tune only BuNN-specific parameters as well as the use of BatchNorm. In Table 5, we report the grid of hyper-parameters that we searched, and denote in bold the best combinations of hyper-parameters. The parameters fixed from Tönshoff et al. [36] are the following:
@@
  • Weight decay 00
  • We use skip connection as implemented in Dwivedi et al. [32] and not in Tönshoff et al. [36]. That is, the skip connection does not skip the non-linearity.
For the BuNN specific parameters, we use $2$ dimensional bundles, whose angles $\theta$ we compute with the help of a small SumGNN architecture using a sum aggregation as defined by $\mathbf{\theta}_{v}^{(\ell)}=\sigma\left(\mathbf{W}_{\text{s}}\mathbf{x}^{(\ell)}_{v}+\mathbf{W}_{\text{n}}\sum_{u\in\mathcal{N}(v)}\mathbf{x}^{(\ell)}_{u}\right)$ where the input dimension is the hidden dimension, the hidden dimension is twice the number of bundles and the output is the number of bundles. The number of SumGNN layers is a hyper-parameter we tune. When it is 00 we use a $2$ layer MLP with hidden dimension also twice number of bundles. For each hyper-parameter configuration, we set the hidden dimension in order to respect to the parameter budget. We use truncated Taylor series for this implementation.
@@
BuNN BuNN-Hop
func struct func struct
Num bundles b $4$, $8$, $16$ $16$ $16$ $16$ $16$
Number of BuNN layers 1 6161\ -\ 61 - 6 $6$ $4$ $6$ $4$
Number of SumGNN layer 0 3030\ -\ 30 - 3 $1$ 00 $6$ $4$
Max expansion degree $K$ $1$  $2$  $4$  $8$  $16$  $32$  $64$  $128$ $128$ $1$ $64$ $64$
Time $t$ $0.1$, $1$, $5$, $6$, $7$, $8$, $9$, $10$, $\infty$ $1$ $6$ N/A N/A
Table 5: Grid of hyper-parameters for peptides-func and peptides-struct.
@@
For the heterophilic graphs we use the source code from Platonov et al. [33] in which we add our layer definition. We report all training parameters that we have tuned. Namely, we use GELU activation functions, the Adam optimizer with learning rate $3\times 10^{-5}$, and train all models for $2000$ epochs and select the best epoch based on the validation set. To compute the bundle maps, we compute the parameters $\theta$ with a GraphSAGE architecture shared across layers ($\phi$ method = shared) or different at each layer ($\phi$ method = not shared), with hidden dimension dimension the number of bundles. The number of layers of this GNN is a hyper-parameter we tuned, which when set to 00 we use a $2$ layer MLP. For each task we manually tuned parameters, which are subsets of the combinations of parameters in the grid from Table 6. The implementation of the heat kernel used is either truncated Taylor series with degree $8$, or the Spectral implementation. We report the best performing combination of parameters in Table 7.
@@
Taylor expansion degree $K$ $N/A$ $N/A$ $8$ $8$ $8$
Time $t$ $100$ $1$ $1$ $1$ $1$
$\phi$ method not shared not shared not shared not shared shared

### v2

@@ Appendix F Experiment details
For peptides-func and peptides-struct we use a fixed parameter budget of $\sim 500k$ as in Dwivedi et al. (2022). We fix hyper-parameters to be the best GCN hyper-parameters from Tönshoff et al. (2023), and tune only BuNN-specific parameters as well as the use of BatchNorm. In Table 5, we report the grid of hyper-parameters that we searched, and denote in bold the best combinations of hyper-parameters. The parameters fixed from Tönshoff et al. (2023) are the following:
@@
  • We use skip connection as implemented in Dwivedi et al. (2022) and not in Tönshoff et al. (2023). That is, the skip connection does not skip the non-linearity.
For the BuNN specific parameters, we use $2$ dimensional bundles, whose angles $\theta$ we compute with the help of a small SumGNN architecture using a sum aggregation as defined by $\mathbf{\theta}_{v}^{(\ell)}=\sigma\left(\mathbf{W}_{\text{s}}\mathbf{x}^{(\ell)}_{v}+\mathbf{W}_{\text{n}}\sum_{u\in\mathcal{N}(v)}\mathbf{x}^{(\ell)}_{u}\right)$ where the input dimension is the hidden dimension, the hidden dimension is twice the number of bundles and the output is the number of bundles. The number of SumGNN layers is a hyper-parameter we tune. When it is 00 we use a $2$ layer MLP with hidden dimension also twice number of bundles. For each hyper-parameter configuration, we set the hidden dimension in order to respect to the parameter budget.
@@
func struct PascalVOC-SP
Num bundles b $4$, $8$, $16$ $16$ $16$ $16$
Number of BuNN layers 1 201201\ -\ 201 - 20 $6$ $4$ $20$
Number of SumGNN layer 0 3030\ -\ 30 - 3 $1$ 00 00
Weight decay 0, 0.1, 0.2, 0.300.10.20.30,\ 0.1,\ 0.2,\ 0.30 , 0.1 , 0.2 , 0.3 $0.2$ $0.2$ $0.2$
Time $t$ $0.1$, $1$, $10$, $100$ $1$ $1$ $1$
Table 5: Grid of hyper-parameters for peptides-func, peptides-struct, and PascalVOC-SP.
@@
For the heterophilic graphs we use the source code from Platonov et al. (2023) in which we add our layer definition. We report all training parameters that we have tuned. Namely, we use GELU activation functions, the Adam optimizer with learning rate $3\times 10^{-5}$, and train all models for $2000$ epochs and select the best epoch based on the validation set. To compute the bundle maps, we compute the parameters $\theta$ with a GraphSAGE architecture shared across layers ($\phi$ method = shared) or different at each layer ($\phi$ method = not shared), with hidden dimension dimension the number of bundles. The number of layers of this GNN is a hyper-parameter we tuned, which when set to 00 we use a $2$ layer MLP. For each task we manually tuned parameters, which are subsets of the combinations of parameters in the grid from Table 6. The implementation of the heat kernel used is either truncated Taylor series with degree $8$, or the spectral implementation. We report the best performing combination of parameters in Table 7. For the NSD baseline we use code from Bodnar et al. (2022) and tune equivalent parameters. We report the grid of hyperparameters in Table 8 and best values in Table 9
@@
Time $t$ $100$ $1.5$ $1$ $1$ $1$
shared $\phi$
@@
Parameters All Values
Hidden dim 64, 128, $64,\ 128,\ 256^{*},512^{*}$
Sheaf dimension $d$ 2, 4, 82482,\ 4,\ 82 , 4 , 8
[…]
Dropout 0.0, 0.20.00.20.0,\ 0.20.0 , 0.2
Learning rate $3\times 10^{-4},3\times 10^{-5}$
Table 8: Parameters searched when tuning NSD on the heterophilic graph benchmark datasets. Parameters marked by ran out of memory on some datasets.
Parameters Best Values
roman-empire amazon-ratings minesweeper tolokers questions
Hidden dim $256$ $256$ $512$ $64$ $64$
Sheaf dim $2$ $4$ $2$ $2$ $2$
Number of layers $8$ $3$ $8$ $5$ $5$
Dropout $0.2$ $0.2$ $0.2$ $0.2$ $0.2$
Learning rate $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
Table 9: Best parameter for NSD on each dataset in the heterophilic graph benchmarks.
Appendix G Empirical runtime
Table 10: Average training time, over $3$ runs, for different architectures of with $5$ layers and hidden dimension of $512$ on the Heterophilic datasets (averaged over $3$ runs). BuNN and NSD use a single $2$ dimensional sheaf. BuNN is implemented spatially. All experiments were performed on an NVIDIA A10 (24GB) GPU.
avg. training time roman-empire amazon-ratings minesweeper tolokers questions
# num nodes 22,662 24,492 10,000 11,758 48,921
# num edges 32,927 93,050 39,402 519,000 153,540
SAGE 1:45 1:43 0:48 1:15 3:01
GAT 2:33 2:42 1:11 3:24 5:22
GT 3:31 4:12 1:52 4:20 7:57
NSD 7:58 9:16 7:13 OOM OOM
BuNN 4:21 5:18 2:48 3:51 9:32
Table 11: Average time per epoch, over $5$ epochs, for different architectures with $6$ layers and $500$k parameters on the LRGB datasets. BuNN is implemented spectrally and uses a single $2$ dimensional bundle. All experiments were performed on an NVIDIA A10 (24GB) GPU.
avg. time / epoch Peptides-func Peptides-struct PascalVOC-SP
# graphs 15,535 15,535 11,355
# nodes 2,344,859 2,344,859 15,955,687
# edges 4,773,974 4,773,974 32,341,644
GCN 5.6s 5.3s 15.2s
GatedGCN 8.5s 8.3s 22.5s
GPS 17.2s 17.2s 38.2s
BuNN 12.7s 12.5s 28.7s
Appendix H Ablation: Positional Encoding ablation
We perform an ablation on the use of PE on the peptides-func and peptides-struct datasets. We retrain a model using the best hyperparameter where we change the used PEs. We compare the two main PE used in graph machin learning, namely Laplacian Positional encoding (LPE) and Random Walk Structural Encodings (RWSE), and we consider using no PE as a baseline. Results can be found in Table 12. We observe that using PE is always beneficial to not using PE, however each task seem to admit a preferred PE, since LPE is better for Peptides-struct and RWSE for Peptides-func.
Table 12: Positional encoding ablation on peptides-func and peptides-struct
RWSE LPE No PE
Peptides-func  Test AP $\uparrow$ $\mathbf{72.76\pm 0.65}$ $72.25\pm 0.51$ $71.76\pm 0.68$
Peptides-struct Test MAE $\downarrow$ $25.02\pm 0.15$ $\mathbf{24.63\pm 0.12}$ $25.32\pm 0.19$
Appendix I Ablation: Importance of $W$ and $b$
We run an ablation on the importance of the $\mathbf{W}$ and $b$ parameters in Equation 2. We retrain a model removing them and compare it to a model trained using them. We report results in Table 13. The results suggest that these parameters help. However, the result suggest that they are not essential to achieve good results, as even without them, the model performs well and beats all but one of the baselines on peptides-func.
Table 13: $W$ and $b$ importance ablation on peptides-func and peptides-struct
with without
Peptides-func  Test AP $\uparrow$ $\mathbf{72.76\pm 0.65}$ $70.75\pm 0.36$
Peptides-struct Test MAE $\downarrow$ $\mathbf{24.63\pm 0.12}$ $25.28\pm 0.32$
Appendix J Tree-NeighborsMatch Task
As an additional synthetic task, we evaluate BuNNs on the Tree-NeighborsMatch task, proposed in Alon & Yahav (2021) to show that MPNNs suffer from over-squashing. We use their code and setup to evaluate the capacity fo BuNNs to alleviate over-squashing. We use their reported results and add our own results, ran using their experimental setup, for a $2$-layer deep BuNN with $t=\infty$ and $32$ bundles. We report the results in Figure 5. We observe that BuNN beat all MPNNs by a large margin, with perfect until $r=6$. As the task gets harder with a larger depth, the accuracy for BuNNs drops slower than the accuracy of MPNNs. These results confirm that BuNNs alleviate over-squashing.
Figure 5: Results for Tree-NeighborsMatch task.